OpenStudy (anonymous):

Given f(x) = 2x2 + 2x and g(x) = x + 1, find the function (f/g)(x).

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

divide the two functions and cancel out the common factors

OpenStudy (anonymous):

whats the excluded value? 0 or -1?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

2x...wen x!=-1

OpenStudy (anonymous):

thanks soo much!

OpenStudy (anonymous):

aagaye?

OpenStudy (anonymous):

u cannot cancel (x+1) wen x=-1 cause then ur doing a 0/0 cancellation and rendering it as 1. but dont worry for the domain of x, for f/g(x) itx 2x.

OpenStudy (anonymous):

f/g (x) = 2(2+x)/(x+1)

OpenStudy (anonymous):

yea,no interesting questions to solve...i'm seeing {H]ouse and helping these guys , parttime ;) :D

OpenStudy (anonymous):

come on chat

OpenStudy (anonymous):

sovle two more and i free to come